OpenStudy (anonymous):

I need help factoring 5x^2-6x+1

OpenStudy (anonymous):

5x^(2)-6x+1 For a polynomial of the form ax^(2)+bx+c, find two factors of a*c (5) that add up to b (-6).In this problem -(1)/(5)*-1=(1)/(5) (which is (c)/(a)) and -(1)/(5)-1=-(6)/(5) (which is ((b)/(a)) , so insert -(1)/(5) as the right hand term of one factor and -1 as the right-hand term of the other factor. (x-(1)/(5))(x-1) Remove the fraction by multiplying the first term of the factor by the denominator of the second term. (5x-1)(x-1)

OpenStudy (anonymous):

Thank you soo much :) I ended up not reading it right.

OpenStudy (anonymous):

your welcome :)

geerky42 (geerky42):

@some_someone

geerky42 (geerky42):

Just saying, don't rustle your jimmies :)

geerky42 (geerky42):

@conami Do you understand and know what to do?